OpenStudy (anonymous):

enter the ordered pair that is the solution to the system of equations graphed below. y=3/2x+1 y=x+2

OpenStudy (mertsj):

Replace y in the second equation with 3/2x + 1 from the first equation. Then solve that equation for x.

OpenStudy (anonymous):

but im not good at fractions can you help me solve it.?

OpenStudy (mertsj):

Write the equation that you got from doing the substitution.

OpenStudy (anonymous):

x+2=3/2x+1

OpenStudy (mertsj):

Subtract x from both sides.

OpenStudy (anonymous):

but idk what that would be like I said im not good at fractions

OpenStudy (mertsj):

3/2 = 1 1/2 Subtract 1 from 1 1/2

OpenStudy (anonymous):

1/2.?

OpenStudy (mertsj):

yes. 3/2x - 1x = 1/2 x

OpenStudy (mertsj):

So now you have: \[2=\frac{1}{2}x+1\]

OpenStudy (mertsj):

Subtract 1 from both sides.

OpenStudy (anonymous):

1=1/2x

OpenStudy (mertsj):

Now multiply both sides by 2

OpenStudy (anonymous):

2=x.?

OpenStudy (anonymous):

(2,4)

OpenStudy (mertsj):

Very good.